If the coefficient of friction between tires and roadway is 0.50, what is the smallest radius at which 

     a car can turn on a horizontal road when its speed is 30 miles/hour? 


1
Expert's answer
2022-06-02T10:26:31-0400

Given:

"\mu=0.5"

"v=30\:\rm mi/hr=13.4\: m/s"


"ma=F"

"\frac{mv^2}{R}=\mu N=\mu m g"

"R=v^2/(\mu g)\\
=(13.4)^2/(0.5*9.8)=37\:\rm m"